calculate the ph, ph=-log(AH+) of a 0.0200 M solution of Ca(NO3)2 to which 5.87 x10^-7 moles of HCL have been added per L of solution. Include ionic strength effects in this calculation and equilibrium constant.
Answer #1

we have

0.0200 M solution of Ca(NO3)2.

Ca(NO3)2 ........> Ca^2+ + 2 NO3-

[Ca^2+] = 0.02 M

and

[NO3-] = 2 * 0.02 = 0.04 M

ionic strength (v) = 1/2 * [0.02 * (+2)^2 + 0.04 * (-1)^2] = 0.06 M

we know,

log (activity coefficient) = - A * Z+ * Z- * sqrt (v)

or

log (activity coefficient) = - 0.51 * 2 * 1 * sqrt (0.06)

or

activity coefficient = 0.5625

thus

HCl is strong acid.

activity of H+ = 5.87 * 10^-7 * 0.5625 = 3.302 * 10^-7 M

pH = -log (activity of H+) = - log (3.302 * 10^-7 ) = 6.48 (answer)